### v1

# Elements of Cohomology for model-theorically finite-dimensional groups and Lie algebras [^1] Keywords: nilpotent group, nilpotent Lie algebra, cohomology, finite-dimensional theory MSC 2020 : 17B56, 03C60

Abstract
We use the cohomology theory for groups invented by Hochshild and Serre to compute the first cohomology group for nilpotent groups that are definable in a finite-dimensional theory. Based on the established cohomological results, we derive some structural results : we prove a weak form of the Frattini argument for definable connected Cartan subgroups and we give a definable version of Maschke’s theorem in the case of a definable abelian $p^{\perp}$-group. The same results hold when one works with Lie algebras instead of groups.
1 Introduction
Under the impetus of Hochschild and Serre (see [7] and [8]), techniques from homological algebra have been successfully applied to the study of modules in the category of groups and of Lie algebras. For groups, cohomological methods play a prominent role in the classification of group extensions and had striking applications in number theory through Galois cohomology (see [10]). In the case of Lie algebras, a number of structural results on finite-dimensional (in the usual linear sense) solvable Lie algebras were obtained by Barnes by using cohomological tools (see [2], [3], and [4]).
Therefore, it would be natural to generalize these methods to the context of groups/Lie algebras and associated modules when one works within the category of definable objects in a given first-order theory. We will consider the case of finite-dimensional theories (precise definitions will be given in the following). The existence of a finite dimension on the collection of interpretable sets, endowed in addition with good algebraic properties generalizing those of linear dimension, will enable us to have better control over constructions arising from homological algebra.
Regarding the scope of our approach to cohomology in this context, we will use two main tools : 1) the exactness of the inflation-restriction sequence for $n=1$, which can be proved by using elementary computations on the cochains 2) the existence of long exact sequences of cohomology groups induced by short exact sequences of $G$-modules. In particular, the Hoshchild-Serre spectral sequence will play no role in our approach.
We formulate our results for groups, but we could have opted for the point of view of Lie algebras (see the Appendix).
The main results we prove are the following :
  • In a finite-dimensional theory, let $G$ be a definable connected nilpotent group and $A$ be a definable connected $G$-module. Suppose that $A^{G}=0$. Then $H^{1}(G,A)=0$.
  • In a finite-dimensional theory, let $G$ be a definable connected nilpotent group and $A$ be a definable connected $G$-module. Suppose that $A^{G}=0$. Then for all definable connected sub-modules $V\leq U$ contained in $A$, one has $(U/V)^{G}=0$.
  • In a finite-dimensional theory, let $G$ be a definable connected group, and let $C$ be a definable connected Cartan subgroup of a definable connected normal subgroup $H$. Suppose that $G/N_{G}(C)$ is abelian. Then
    $G=HN_{G}(C).$
  • In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module $p$-elementary as an abelian group. Suppose that $A^{T}=0$. Then $A=A_{1}\oplus\dots\oplus A_{r}$, where each $A_{i}$ is a definable connected $T$-irreducible sub-module of $A$.
In the second section, we recall some basic results about the cohomology of groups. In the third section, we consider the case of a finite nilpotent group $G$ and a finite $G$-module and we give an elementary proof of the vanishing of the first cohomology group when the submodule of invariants is trivial. In the fourth section, we generalize the cohomological results for finite groups to the case of nilpotent groups definable in a finite-dimensional theory. In the fifth section, based on the established cohomological results, we derive a weak form of the Frattini argument for connected definable Cartan subgroups. In the sixth section, we prove a definable Maschke’s theorem for definable abelian groups without $p$-torsion acting on definable abelian $p$-elementary groups. Finally, in an appendix, we explain what happens to our results when we consider Lie algebras instead of groups.
We thank Adrien Deloro for his many comments and suggestions.
2 Preliminaries on the cohomology of groups
We recall some basic facts regarding the cohomology of groups; for the sake of completeness, we give proofs when it is needed. For general references on this topic, we refer to [7] and to [10].
Definition 2.1.
A $G$-module is the data of ($G$,AA), where $G$ is a group and $A$ is an abelian group, equipped with a group action $\cdot"$ of $G$ on $A$. Equivalently, this amounts to giving a group homomorphism $\rho$ between $G$ and $\operatorname{Aut}(A)\subseteq\operatorname{End}(A)$.
Definition 2.2.
Let $(G,A)$ be a $G$-module. We define a cochain complex as follows :
  • $C^{0}(G,A)=A$
  • For every integer $n\geq 1$, $C^{n}(G,A)$ denotes the abelian group of functions between $G^{n}$ and AA
The differential $d_{n}$ between $C^{n}(G,A)$ and $C^{n+1}(G,A)$ is defined, for every map $f$ in $C^{n}(G,A)$, by:
$d_{n}f(g_{1},\dots,g_{n+1})$
$=g_{1}\cdot f(g_{2},\dots,g_{n+1})+(-1)^{n+1}f(g_{1},\dots,g_{n})$
$+\sum_{1\leq s\leq n}(-1)f(g_{1},\dots,g_{s}g_{s+1},\dots,g_{n+1})$
We have $d_{n+1}\circ d_{n}=0$. We set $C(G,A)=(C^{n}(G,A))_{n\in\mathbb{N}}$.
We may now define the associated cohomology groups :
Definition 2.3.
Given the complex defined above, we introduce for every integer $n$ the $n$-th cohomology group, denoted by $H^{n}(G,A)$ :
$H^{n}(G,A)=\ker(d_{n})/\operatorname{im}(d_{n-1}).$
The group $G$ acts on the cochain groups of the form $C^{n}(H,A)$, for $H$ a normal subgroup, in a way that is compatible with the differentials : let $f\in C^{n}(H,A)$ and $x\in G$, then $(x\cdot f)(h_{1},\dots,h_{n})=x\cdot f(h_{1}^{x},\dots,h_{n}^{x})$.
For $n=0$ and $n=1$, it is possible to give a convenient algebraic description of the cohomology groups.
Fact 2.4.
  1. 1.
    $H^{0}(G,A)=A^{G}=\{a\in A:g\cdot a=a,\text{for all $g\in G$}\}$.
  2. 2.
    $H^{1}(G,A)=\operatorname{Der}(G,A)/\operatorname{IDer}(G,A)$, where
    $\operatorname{Der}(G,A)=\{f\in C^{1}(G,A):f(xy)=x\cdot f(y)+f(x)\}$
    and
    $\operatorname{IDer}(G,A)=\{f\in C^{1}(G,A):f(x)=x\cdot a-a\penalty 10000\ \text{for some $a$ in $A$}\}.$
    In particular, if $f\in\operatorname{Der}(G,A)$, then $f(1)=0$.
Remark 2.5.
We observe that $G$ acts trivially on $H^{1}(G,A)$: let $f]\in H^{1}(G,A)$, where $f$ is a derivation, and let $x\in G$; then $(x\cdot f)(y)-f(y)=x\cdot f(y^{x})-f(y)=x\cdot(x^{-1}\cdot f(yx)+f(x^{-1}))-f(y)=f(yx)+f(1)-f(x)-f(y)=y\cdot f(x)+f(y)-f(x)-f(y)$ for all $y\in G$; in other words, $(x\cdot f)-f$ is an inner derivation and thus $x\cdot f]=[f]$.
We now turn to the question of the relations between exact sequences of $G$-modules and cohomology.
Fact 2.6.
Every short exact sequence of $G$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces a short exact sequence of complexes $0\rightarrow C(G,A)\overset{i^{*}}{\rightarrow}C(G,B)\overset{p^{*}}{\rightarrow}C(G,C)\rightarrow 0$, where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$.
Proof.
Since $i$ and $p$ commute with the action of $G$, the formula for the differential allows us to show that $i^{*},p^{*}$ commute with the differential. Fix an integer $n$.
Let us show that $i^{*}$ is injective. Indeed, for $\phi\in C^{n}(G,A)$, $i^{*}(\phi)=0$ iff $i(\phi(\overline{x}))=0$ for all $\overline{x}\in G^{n}$ iff $\phi(\overline{x})=0$ for all $\overline{x}\in G^{n}$ by injectivity of $i$.
Now, we turn to the equality $\operatorname{im}(i^{*})=\ker(p^{*})$. Let $\phi\in\operatorname{im}(i^{*})\subseteq C^{n}(G,B)$; then there exists $\psi\in C^{n}(G,A)$, such that $\phi=i\circ\psi$ and hence $p^{*}(\phi)=p\circ i\circ\psi$. Consequently, for every $\overline{x}\in G^{n}$, $p(i(\psi(\overline{x}))=0$ since $\operatorname{im}(i)\subseteq\ker(p)$, and thus $\phi\in\ker(p^{*})$.
Conversely, let $\phi\in\ker(p^{*})$; then $p(\phi(\overline{x}))=0$ for every $\overline{x}\in G^{n}$. Thus $\phi(\overline{x})\in\operatorname{im}(i)$, since $\operatorname{im}(i)=\ker(p)$. Therefore for every $\overline{x}$ there exists an unique $y_{\overline{x}}\in A$, such that $i(y_{\overline{x}})=\phi(\overline{x})$ by injectivity of $i$. We can define an element of $C^{n}(G,A)$ by $\overline{x}\mapsto y_{\overline{x}}$.
Now, we prove the surjectivity of $p^{*}$. Let $s:C\rightarrow B$ be a set-theoretic section, i.e., $p\circ s=Id_{C}$. Let $\phi\in C^{n}(G,C)$; for $\overline{x}\in G$, consider $s(\phi(\overline{x}))$. Therefore, it suffices to define a cochain $\psi\in C^{n}(G,B)$ by $\psi(\overline{x})=s(\phi(\overline{x}))$ to get a map such that $p^{*}(\psi)=\phi$. ∎
Recall now an important property of the sequence of cohomology groups of a complex: to a short exact sequence of complexes, one can canonically associate a long exact sequence of cohomology groups.
Fact 2.7.
[14, Theorem 1.3.1], [9, Proposition 6.9]
Let $0\rightarrow C^{\prime}\rightarrow C\rightarrow C^{\prime\prime}\rightarrow 0$ be a short exact sequence of cochain complexes. Then there exists a long exact sequence of cohomology groups of the following form :
$\dots\rightarrow H^{n}(C^{\prime})\rightarrow H^{n}(C)\rightarrow H^{n}(C^{\prime\prime})$
$\overset{\delta_{n}}{\rightarrow}H^{n+1}(C^{\prime})\rightarrow H^{n+1}(C)\rightarrow\dots$
The morphisms $(\delta_{n})_{n\in\mathbb{Z}}$ are called connecting morphisms.
As a consequence, an exact sequence of $G$-modules induces a long exact sequence of cohomology groups.
We now introduce the Hochschild-Serre inflation-restriction sequence.
Definition 2.8.
Let $(G,A)$ be a $G$-module and let $H$ be a normal subgroup of $G$.
  • For every $n\in\mathbb{N}$, we define the restriction map $\operatorname{res}:C^{n}(G,A)\longrightarrow C^{n}(H,A),f\mapsto f|_{H}$.
  • We define the inflation map $\inf:C^{n}(G/H,A^{H})\longrightarrow C^{n}(G,A),f\mapsto f\circ\pi$, where $\pi$ denotes the canonical projection between $G$ and $G/H$.
Since the inflation and restriction morphisms commute with the differentials, we obtain a sequence of complexes. In particular, we deduce a sequence $H^{1}(G/H,A^{H})\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)$. In fact, this sequence is exact, and it can be proved by an elementary algebraic argument.
Fact 2.9.
[10, Proposition 4, §6, chap. VII])
The sequence $0\rightarrow H^{1}(G/H,A^{H})\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)$ is exact.
Proof.
  • $\inf$ is injective : let $f\in\operatorname{Der}(G,A)$, such that $f$ is constant on classes modulo $H$ and such that for all $g\in G$, $f(g)=g\cdot a-a$ for some $a\in A$. Then for all $h\in H$, $0=f(1)=f(h)=h\cdot a-a$, so $h\cdot a=a$.
  • $\operatorname{im}(\inf)\subseteq\ker(\operatorname{res})$ : let $f\in\operatorname{Der}(G,A)$ be constant on classes modulo $H$; we have $f(1)=0=f(h)$ for all $h\in H$.
  • $\ker(\operatorname{res})\subseteq\operatorname{im}(\inf)$. Let $f]\in H^{1}(G,A)$, such that $f|_{H}]=[0]$ in $H^{1}(H,A)$; thus there exists $a\in A$ such that $f(h)=h\cdot a-a$ for all $h\in H$. Up to considering the $f$-cohomologue $G$-derivation $\hat{f}$ defined by $\hat{f}(g)=f(g)-(g\cdot a-a)$, we may suppose that $f|_{H}=0$. Hence $f(gh)=f(g)+g\cdot f(h)=f(g)$ and $f$ is constant on classes modulo $H$; so $f(g)=f(gh)=f(hg)=f(h)+h\cdot f(g)=h\cdot f(g)$ since $H$ is a normal subgroup.
We can refine the image of the restriction morphism.
Fact 2.10.
$H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)^{G/H}$.
Proof.
Let $\phi]\in H^{1}(H,A)$ such that $\phi]=[\psi|_{H}]$ for $\psi\in\operatorname{Der}(G,A)$. As $\psi$ is a $G$-derivation, we are done by the same calculations as of Remark 2.5. ∎
The following Lemma will be very useful in the following.
Lemma 2.11.
Let $(G,A)$ be a $G$-module and let $H$ be a central subgroup of $G$, such that $H\leq C_{G}(A)$. Suppose that $A^{G}=0$. Then $H^{1}(H,A)^{G/H}=0$ and $H^{1}(G/H,A)\simeq H^{1}(G,A)$.
Proof.
Since $A^{H}=A$, by fact 2.9 and fact 2.10, the sequence $0\rightarrow H^{1}(G/H,A)\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)^{G/H}$ is exact. Therefore, it suffices to prove that
$H^{1}(H,A)^{G/H}=0.$
Let $f\in\operatorname{Der}(H,A)$; suppose that for all $x\in G$, $x\cdot f]=[f]$, i.e., for all $x\in G$ and for all $h\in H$, we have $x\cdot f(h^{x})-f(h)=x\cdot f(h)-f(h)=h\cdot a_{x}-a_{x}=0$ for some $a_{x}\in A$, since $H$ acts trivially on $A$ and $H\leq Z(G)$. In other words, for every $x\in G$, the equality $x\cdot f(h)=f(h)$ holds, and thus $f(h)\in A^{G}=0$; finally, we obtain that $f=0$. ∎
3 The case of finite groups
In this section, we give a quite elementary proof of the vanishing of the first cohomology group when $A^{G}=0$ for a finite nilpotent group $G$ and a finite $G$-module AA. This seems to be a quite classical result in the cohomology of finite groups, but our proof does not make use of Tate’s cohomology and it is more suitable to generalization to a model-theoretic framework. We draw inspiration of the proof strategy used by Barnes in [3] for finite-dimensional Lie algebras.
Let us recall a basic but fundamental lemma in representation theory : Schur’s Lemma.
Fact 3.1.
Let $A$ be a simple $R$-module, i.e $A$ has no non-trivial proper submodule. Then $C_{\operatorname{End}(A)}(R)$ is a skew-field.
We begin by the abelian case.
Lemma 3.2.
Let $G$ be a finite abelian group and $A$ be a finite $G$-module. Suppose that $A^{G}=0$. Then $H^{1}(G,A)=0$.
Proof.
We proceed by induction on the length, $\lg$, of $A$ in the category of finite $G$-modules. Since $A$ is finite, the length is also finite. If $A=0$, we are done.
  • $\lg(A)=1$, i.e., $A$ is irreducible. First, we way suppose that $G$ acts faithfully by Lemma 2.11 and so we have an embedding $\rho:G\hookrightarrow\operatorname{End}(A)$. Consider $R$ the subring of $\operatorname{End}(A)$ generated by $G$ and notice that $A$ is a simple $R$-module. By Schur’s Lemma, $C_{\operatorname{End}(A)}(R)$ is a skew field containing $\rho(G)$; this is even a commutative field by Wedderburn’s theorem.
    Now, let $f:G\rightarrow A$ be a non-trivial $G$-derivation. Let $y\in G-\{1\}$, such that $f(y)\neq 0$; then $(\rho(y)-Id)$ is invertible. Let $a=(\rho(y)-Id)^{-1}\cdot f(y)$. For $x\in G$, by commutativity of $G$ and the definition of a derivation, we have $(\rho(y)-Id)\cdot f(x)=\rho(x)\cdot f(y)-f(y)$, thus, $f(x)=(\rho(y)-Id)^{-1}\cdot\rho(x)\cdot f(y)-(\rho(y)-Id)^{-1}\cdot f(y)=\rho(x)\cdot a-a$ by commutativity again. This shows that $H^{1}(G,A)=0$.
  • $\lg(A)>1$. Then, there exists a proper non-trivial $G$-module $B$, such that $\lg(B)<\lg(A)$ and $\lg(A/B)<\lg(A)$. Note that by hypothesis $B^{G}=0$. By induction on the length, we may therefore assume that $H^{1}(G,B)=0$. The long exact cohomology sequence will now come into play. Indeed, we have a short exact sequence of $G$-modules $0\rightarrow B\rightarrow A\rightarrow A/B\rightarrow 0$ which, by virtue of Fact 2.7, induces the following long exact cohomology sequence :
    $0\rightarrow B^{G}\rightarrow A^{G}\rightarrow(A/B)^{G}$
    $\overset{\delta_{0}}{\rightarrow}H^{1}(G,B)\rightarrow H^{1}(G,A)\rightarrow H^{1}(G,A/B)\rightarrow\dots$
    By hypothesis, $A^{G}=B^{G}=0$ and by induction hypothesis, $H^{1}(G,B)=0$, so the exactness of the sequence then gives us $(A/B)^{G}=0$; we may therefore apply once more the induction hypothesis and use the exactness of the long cohomology sequence to establish $H^{1}(G,A)=0$.
Corollary 3.3.
Let $G$ be a finite nilpotent group and $A$ be a finite $G$-module. Suppose that $A^{G}=0$ and $A$ is $G$-irreducible. Then $H^{1}(G,A)=0$.
Proof.
We proceed by induction on the nilpotency class of $G$. We may suppose that $G$ is not abelian; since $G$ is nilpotent, we have $Z(G)\neq 1$. Notice that $A^{Z(G)}$ is a $G$-module. By irreducibility of $A$, we are reduced to two cases.
  • $A^{Z(G)}=A$. Lemma 2.11 gives $H^{1}(Z(G),A)^{G/Z(G)}=0$; so, the inflation-restriction exact sequence entails $H^{1}(G/Z(G),A)\simeq H^{1}(G,A)$ and we are done by induction.
  • $A^{Z(G)}=0$. Since $H^{1}(Z(G),A)=0$ (abelian case) and $H^{1}(G/Z(G),0)=0$, we have $H^{1}(G,A)=0$ by considering again the inflation-restriction exact sequence.
Corollary 3.4.
Let $G$ be a finite nilpotent group and $A$ be a finite $G$-module. Suppose that $A^{G}=0$. Then $H^{1}(G,A)=0$.
Proof.
We proceed exactly in the same way as in the proof of Lemma 3.2. ∎
To extend the proof to the category of definable $G$-modules in a first order theory, we need two main ingredients : 1) a notion of irreducible module that allows inductive reasoning on the lenght 2) an appropriate version of Schur’s Lemma regarding this irreducible category. In the next section, we will see that finite-dimensional theories satisfy these requirements.
4 Finite-dimensional nilpotent groups
We will study group cohomology inside the framework of finite-dimensional theories.
Definition 4.1.
[13, Definition 1.1] A theory $T$ is said to be finite-dimensional if there exists a dimension function, $\dim$, defined on the collection of all interpretable sets in models of $T$, with values in $\mathbb{N}$, such that for a formula $\phi(x,y)$ and interpretable sets $X,Y$, the following properties are satisfied :
  • (Invariance) If $a\equiv a^{\prime}$, then $\dim(\phi(x,a))=\dim(\phi(x,a^{\prime}))$.
  • (Finiteness) $\dim(X)=0$ iff $X$ is finite.
  • (Union) $\dim(X\cup Y)=\max\{\dim(X),\dim(Y)\}$.
  • (Fibration) If $f:X\longrightarrow Y$ is an interpretable map, such that $\dim(f^{-1}(y))\geq d$ for all $y\in Y$, then $\dim(X)\geq\dim(Y)+d$.
The dimension extends to type-definable sets, i.e., bounded intersections of definable sets, by taking the infimum of the dimensions of the definable sets which appear in the intersection.
A strongly minimal theory, the theory of a group of finite Morley rank and of an o-minimal structure are examples of finite-dimensional theories (for o-minimal structures, we must restrict to the collection of definable sets and interpretable group quotients). This model-theoretic framework is suitable for inductive reasoning when working with groups.
Fact 4.2.
[13, Proposition 2.3] Let $H\leq G$ be two type-definable groups in a finite-dimensional theory. If $H$ has infinite index in $G$, then $\dim(H)<\dim(G)$.
A definable group is said to be connected if it has no definable subgroup of finite index.
Another remarkable property of finite-dimensional theories concerns the possibility of obtaining a definable version of Schur’s Lemma.
Fact 4.3.
(adapted from [5]) Let $T$ be a finite-dimensional theory. Let $V$ be a definable connected abelian group and $S\leq\operatorname{End}_{def}(V)$ an invariant, i.e., a bounded union of type-definable sets, and unbounded subring of definable endomorphisms. Suppose that $V$ is an irreducible $S$-module, i.e., every proper definable connected $S$-invariant subgroup is trivial. Then $T=C_{End_{def}(V)}(S)$ is a definable skew-field.
In the following, unless explicitly stated otherwise, an irreducible module will be understood in the sense indicated above, i.e., in reference to the definable connected category.
Lemma 4.4.
In a finite-dimensional theory, let $(G,A)$ be a definable $G$-module. Suppose that $G$ is connected and abelian, that $A$ is connected and that $A^{G}=0$. Then $H^{1}(G,A)=0$.
Proof.
First, according to Corollary 2.11, up to considering $G/C_{G}(A)$, we may assume that $G$ acts faithfully. The result is trivial for $A=0$. We proceed by induction on the length, $\lg$, of $A$ as a $G$-module (in the definable connected category).
  • First, suppose $\lg(A)=1$, i.e., $A$ is irreducible. We denote by $\rho:G\rightarrow\operatorname{Aut}_{def}(A)\subseteq\operatorname{End}_{def}(A)$ the action. By Fact 4.3, we can linearize : there exists a definable skew-field $\mathbb{K}$ such that $\rho(G)\leq\mathbb{K}$. Up to considering $Z(\mathbb{K})$, which is also definable infinite, we may assume that $\mathbb{K}$ is commutative. Now, the proof goes as in the Lemma 3.2.
  • We may therefore assume $\lg(A)>1$; then there exists a non-trivial definable connected $G$-module $B$ strictly contained in $A$ such that $\lg(B)<\lg(A)$ and $\lg(A/B)<\lg(A)$. Now, the proof goes as in the Lemma 3.2.
At this point, we must consider a subclass of finite-dimensional theories.
Definition 4.5.
A finite-dimensional theory is said to be finite-dimensional if it satisfies the following property :
Any definable group admits a connected component, i.e., a smallest definable subgroup of finite index.
A strongly minimal theory, the theory of an o-minimal structure, the theory of a group of finite Morley rank belong to this subclass. In a finite-dimensional theory, a definable irreducible $G$-module has no proper infinite definable submodule. Moreover, an infinite definable nilpotent group has infinite center (see [1, Lemma 5.1, I]).
Theorem 4.6.
In a finite-dimensional theory, let $(G,A)$ be a definable $G$-module. Suppose that $G$ is connected and nilpotent, that $A$ is connected, and that $A^{G}=0$. Then $H^{1}(G,A)=0$.
Proof.
We proceed by induction on the nilpotency class of $G$. By Lemma 4.4, we may suppose that $G$ is not abelian. Notice that $G/Z(G)^{\circ}$ and $Z(G)^{\circ}$ are infinite definable connected nilpotent groups of nilpotency class strictly inferior.
First, suppose that $A$ is irreducible (in the definable connected category) and consider the definable submodule $A^{Z(G)^{\circ}}$. Then we have two cases :
  • $A^{Z(G)^{\circ}}=A$. By Lemma 2.11, $H^{1}(G,A)\simeq H^{1}(G/Z(G)^{\circ},A)$, and we are done by induction on the nilpotency class.
  • $A^{Z(G)^{\circ}}$ is finite. In fact, $A^{Z(G)^{\circ}}=0$ since $A^{G}=0$ and $G$ is connected. Notice that $H^{1}(G/Z(G)^{\circ},A^{Z(G)^{\circ}})=0=H^{1}(Z(G)^{\circ},A)$ by Lemma 4.4 and we are done by fact 2.9.
Now, we proceed by induction on the length of $A$ as a $G$-module exactly in the same way as in the proof of Lemma 4.4. ∎
We conclude this section with a corollary on composition factors.
Corollary 4.7.
In a finite-dimensional theory, let $G$ be a definable connected nilpotent group and let $A$ be a definable connected $G$-module. Suppose that $A^{G}=0$. Then for all definable connected sub-modules $V\leq U$ contained in $A$, we have $(U/V)^{G}=0$.
Proof.
Indeed, we observe that $U^{G}=0=V^{G}$; then it suffices to consider the long exact cohomology sequence :
$0\rightarrow V^{G}\rightarrow U^{G}\rightarrow(U/V)^{G}$
$\overset{\delta_{0}}{\rightarrow}H^{1}(G,V)\dots$
Theorem 4.6 allows us to conclude. ∎
5 A weak Frattini argument for definable connected Cartan subgroups
In the theory of finite groups, $p$-Sylow subgroups play a fundamental role and often allow to obtain structural results about the ambient group. In particular, the Frattini argument states that, for a finite group $G$, a normal subgroup $H$, and a $p$-Sylow subgroup $S$ of $H$, we have the following :
$G=HN_{G}(S),$
where $N_{G}(S)$ denotes the normalizer of $S$ in $G$. In fact, the image of $S$ under an inner automorphism $inn(g)$ remains a $p$-Sylow subgroup of $H$; by the conjugacy of $p$-Sylow subgroups in $H$, there exists $h\in H$, such that $S^{h}=S^{g}$, that is, such that $h^{-1}g\in N_{G}(S)$.
Outside the world of finite groups, Cartan subgroups, i.e., nilpotent self-normalizing subgroups can be seen as a good proxy for $p$-Sylow subgroups. In general, the existence and the conjugation of Cartan subgroups are not guaranteed. Nevertheless, in a finite-dimensional theory, if there exist definable connected Cartan subgroups, the conclusion of the Frattini argument may still be true without any conjugation assumption. We follow the cohomological treatment for finite-dimensional Lie algebras given by Barnes in [4].
Definition 5.1.
In a finite-dimensional theory, let $G$ be a definable connected group and let $C$ be a definable connected nilpotent subgroup such that
$N_{G}(C)=C.$
We then say that $C$ is a definable connected Cartan subgroup.
Proposition 5.2.
In a finite-dimensional theory, let $G$ be a definable connected group, and let $C$ be a definable connected Cartan subgroup of a definable connected normal subgroup $H$. Suppose that $G/N_{G}(C)$ is abelian. Then
$G=HN_{G}(C).$
Proof.
We set $N=N_{G}(C)$ and we consider the definable connected $C$-module $A=G/N$.
Let us prove that $A^{C}=0$. Indeed, let $\overline{a}\in A$ be such that $c\cdot\overline{a}=\overline{a}$ for all $c\in C$; then $a^{-1}a^{c}\in H\cap N$. But $H$ is normal subgroup and $N_{H}(C)=C$, so $a^{-1}a^{c}\in C$ for all $c$, hence $a\in N_{G}(C)$.
Now, consider the $C$-module
$(G/N)/((HN)/N)\simeq G/(HN).$
If this module were nontrivial, Corollary 4.7 would imply that $C$ acts non-trivially, which contradicts the fact that $H$ contains $C$. Therefore, we conclude that
$G=HN.\qed$
6 Definable Maschke’s theorem
We will provide another application of the cohomological study of model-theorically finite-dimensional nilpotent groups by proving a definable version of Maschke’s theorem. We will closely follow the argument for the case of ranked groups given by Tindzhogo Ntsiri in [11]. We will need the following definability lemma :
Fact 6.1.
([13, Proposition 3.6], extracted from [5]) In a finite-dimensional theory, let $\mathbb{K}$ be a skew-field such that :
  • there is an upper bound on dimensions of type-definable subsets of $\mathbb{K}$.
  • $\mathbb{K}$ contains an invariant unbounded subset.
Then $\mathbb{K}$ is definable.
Proposition 6.2.
In a finite-dimensional theory, let $V$ be a definable connected $p$-elementary abelian group and let $R$ be an invariant commutative ring of definable endomorphisms of $V$ generated by an infinite definable set $X$. Suppose that $R^{p}=R$. Let $W$ be a definable connected irreducible (in the definable category) sub-$R$-module, such that $V/W$ is also $R$-irreducible (in the definable category). Then $W$ admits a definable $R$-invariant complement.
Proof.
First, notice that $R$ is unbounded. By considering the action of $R$ on $W$ and $V/W$ and by applying Fact 4.3, the rings $C_{\operatorname{End}_{def}(W)}(R)=K_{W}$ and $C_{\operatorname{End}_{def}(V/W)}(R)=K_{V/W}$ are in fact definable skew-fields. Therefore $\operatorname{ann}_{R}(W)$ and $\operatorname{ann}_{R}(V/W)$ are prime ideals.
We may suppose that $\operatorname{ann}_{R}(W)=\operatorname{ann}_{R}(V/W)$. Indeed, let $0\neq r\in\operatorname{ann}_{R}(W)-\operatorname{ann}(V/W)$. The definable subgroups $\operatorname{im}(r)$ and $W\subseteq\ker(r)$ are $R$-invariant since $R$ is commutative. Then $\ker(r)/W$ and $(\operatorname{im}(r)+W)/W$ are $R$-invariant, and by irreducibility of $V/W$, we have $V=\operatorname{im}(r)+W$. Moreover, $\dim(V)=\dim(\operatorname{im}(r))+\dim(\ker(r))=\dim(\operatorname{im}(r))+\dim(\ker(r))-\dim(W\cap\operatorname{im}(r))$; thus $\dim(W\cap\operatorname{im}(r))=0$ and $\operatorname{im}(r)\cap W$ is trivial by irreducibility of $W$.
Conversely, let $0\neq r\in\operatorname{ann}_{R}(V/W)-\operatorname{ann}_{R}(W)$. Then $\operatorname{im}(r)=W$ and $\ker(r)\cap W$ is trivial by irreducibility of $W$. Therefore $V=W\oplus\ker(r)$.
Now, let $x\in\operatorname{ann}_{R}(W)=\operatorname{ann}_{R}(V/W)=I$; since $R=R^{p}$, there exists $y\in R$ such that $x=y^{p}$. In the domain $R/I$, we have $\overline{y}^{p}=\overline{x}=0$, so $y\in I$. Then $y\cdot V\subseteq W$ and $y^{2}\cdot V=\{0\}$. Finally, $x=0$ and $R$ embeds in $K_{W}$ and in $K_{V/W}$; by considering $Z(K_{W})$ (respectively, $Z(K_{V/W})$), we may suppose that $K_{W}$ (respectively, $K_{V/W}$) is commutative. Therefore $R$ acts by (scalar)-automorphisms on $W$ and $V/W$.
Now, we may suppose that $R$ acts by automorphisms on $V$ itself. Indeed, let $0\neq r\in R$ and suppose $\ker(r)\neq 0$. Then by irreducibility of $W$, $\ker(r)\cap W=0$; so by irreducibility of $V/W$, $V=W\oplus\ker(r)$.
In $\operatorname{End}_{def}(V)$, we consider the subring $K=R\cdot R^{-1}$; this is a commutative invariant and unbounded field. Let $Z\subseteq K$ be a type-definable subset. For $0\neq v\in V$, the evaluation map $r\mapsto r\cdot v$ is injective since $K$ acts by automorphisms; thus $\dim(Z)=\dim(Z\cdot v)\leq\dim(V)$. By Lemma 6.1, $K$ is in fact a definable field. As a consequence, $W$ is a $K$-subspace of finite linear dimension and admits a definable $R$-invariant complement. ∎
Corollary 6.3.
(compare with [11, Proposition 2.13]) In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module that is $p$-elementary as an abelian group. Suppose that $A^{T}=0$. Then every definable connected irreducible sub-module of $A$ admits a definable $T$-invariant complement.
Proof.
By Corollary 4.7, the hypothesis $V^{T}=0$ is still true for definable connected $T$-invariant subgroups and sections of $V$. We proceed by induction on the length of $V$ as a definable connected $T$-module. Suppose $V\neq 0$; let $W$ be a proper definable connected irreducible sub-module. By induction, we may assume that $0\neq V/W$ is irreducible.
Let $R$ be the subring of $\operatorname{End}_{def}(V)$ generated by $T$. Since $T$ is abelian and $T^{p}=T$, we have $R^{p}=R$. Moreover, $W$ and $V/W$ are $R$-irreducible (in the definable category) : let $W_{1}$ be a finite $R$-sub-module of $W$; in particular, this is a $T$-module, but, by connectedness of $T$, the action is trivial and $W_{1}=0$ since $V^{T}=0$; the proof for $V/W$ is similar since $(V/W)^{T}=0$. Then, we are done by Proposition 6.2. ∎
Corollary 6.4.
In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module that is $p$-elementary as an abelian group. Suppose that $A^{T}=0$. Then $V=V_{1}\oplus\dots\oplus V_{r}$, where each $V_{i}$ is a definable connected $T$-irreducible module.
Proof.
We may suppose that $V$ is not irreducible; let $V_{1}$ be a proper non-trivial definable connected sub-module. We proceed by induction on the length of $V_{1}$. If $\lg(V_{1})=1$, we are done by corollary 6.3. We may suppose that $\lg(V_{1})>1$; so, there exists a proper non-trivial definable connected submodule $V_{2}$ such that $\lg(V_{2})<\lg(V_{1})$ and $\lg(V_{2}/V_{1})<\lg(V_{1})$, and we can conclude by induction. ∎
7 Appendix : Lie algebras
This appendix deals with the case of Lie algebras. Indeed, in [8], Hochshild and Serre defines a cohomology theory for Lie algebras that parallels the theory for groups.
Definition 7.1.
A $\mathfrak{g}$-module is the data of ($\mathfrak{g}$,AA), where $\mathfrak{g}$ is a Lie algebra over a field $K$ and $A$ a $K$-vector space, equipped with a Lie algebra action $\cdot"$ of $\mathfrak{g}$ on $A$. Equivalently, this amounts to giving a Lie algebra homomorphism $\rho$ between $\mathfrak{g}$ and $\operatorname{End}_{K}(A)$.
Definition 7.2.
Let $(\mathfrak{g},A)$ be a $\mathfrak{g}$-module. We define a cochain complex as follows :
  • $C^{0}(\mathfrak{g},A)=A$
  • For every integer $n\geq 1$, $C^{n}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g}^{n},A):f\text{ is alternating}\}$, where $\operatorname{Hom}(\mathfrak{g}^{n},A)$ denotes the set of applications between $\mathfrak{g}^{n}$ and $A$, which are $K$-linear in each variable (we say also $K$-multilinear). Recall that a $K$-multilinear application $f$ is alternating if $f(x_{1},\dots,x_{n})=0$ whenever two arguments among the $x_{i}$ are equal.
The differential $d_{n}$ between $C^{n}(\mathfrak{g},A)$ and $C^{n+1}(\mathfrak{g},A)$ is defined, for every map $f$ in $C^{n}(\mathfrak{g},A)$, by :
$d_{n}f(g_{1},\dots,g_{n+1})$
$=\sum_{1\leq s<t\leq n+1}(-1)^{s+t-1}f([g_{s},g_{t}],g_{1},\dots,\hat{g_{s}},\hat{g_{t}},\dots,g_{n+1})$
$+\sum_{1\leq s\leq n+1}(-1)^{s}g_{s}\cdot f(g_{1},\dots,\hat{g_{s}},\dots,g_{n+1})$
The Lie algebra $\mathfrak{g}$ acts on the cochain groups in a way that is compatible with the differentials :
Let $f\in C^{n}(\mathfrak{g},A)$ and $x\in\mathfrak{g}$, we define $x\cdot f$ by the formula
$(x\cdot f)(y_{1},\dots,y_{n})=x\cdot f(y_{1},\dots,y_{n})-\sum_{1\leq i\leq n}f(y_{1},\dots,y_{i-1},[x,y_{i}],y_{i+1},\dots,y_{n}),$
for all $y_{1},\dots,y_{n}\in\mathfrak{g}$. This action of $\mathfrak{g}$ is still valid if we replace $\mathfrak{g}$ by an ideal $\mathfrak{h}$ in the cochain groups.
For $n=0$ and $n=1$, it is also possible to give a convenient algebraic description of the cohomology groups.
Fact 7.3.
  1. 1.
    $H^{0}(\mathfrak{g},A)=A^{\mathfrak{g}}=\{a\in A:g\cdot a=0,\text{for all $g\in\mathfrak{g}$}\}$.
  2. 2.
    $H^{1}(\mathfrak{g},A)=\operatorname{Der}(\mathfrak{g},A)/\operatorname{IDer}(\mathfrak{g},A)$, where
    $\operatorname{Der}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g},A):f([x,y])=x.f(y)-y.f(x)\}$
    and
    $\operatorname{IDer}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g},A):f(x)=x.a\penalty 10000\ \text{for some $a$ in $A$}\}.$
All the others statements of section 2 are still true when passing to Lie algebras, and the proofs require minor adaptations. In particular, we have the following :
Fact 7.4.
  • Every short exact sequence of $\mathfrak{g}$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces a short exact sequence of complexes $0\rightarrow C(\mathfrak{g},A)\overset{i^{*}}{\rightarrow}C(\mathfrak{g},B)\overset{p^{*}}{\rightarrow}C(\mathfrak{g},C)\rightarrow 0$, where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$.
  • The sequence $0\rightarrow H^{1}(\mathfrak{g}/\mathfrak{h},A^{\mathfrak{h}})\overset{\inf}{\rightarrow}H^{1}(\mathfrak{g},A)\overset{\operatorname{res}}{\rightarrow}H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}$ is exact.
Remark 7.5.
As pointed out by Moreno Invitti, we can obtain the first six terms of the long exact sequence of cohomology groups induced by a short exact sequence of $\mathfrak{g}$-modules when $\mathfrak{g}$ is an abstract Lie ring. Indeed, it suffices to apply the Snake Lemma to the commutative diagram whose lines are $A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ and $0\rightarrow\operatorname{Der}(\mathfrak{g},A)\overset{i^{*}}{\rightarrow}\operatorname{Der}(\mathfrak{g},B)\overset{p^{*}}{\rightarrow}\operatorname{Der}(\mathfrak{g},C)$. Since we do not use the long exact sequence of cohomology groups for higher values of $n$, all the results of this appendix are still true for Lie rings that are definable in a finite-dimensional theory. Some generalizations of the results of this paper are currently under investigation by Moreno Invitti.
Lemma 7.6.
Let $(\mathfrak{g},A)$ be a $\mathfrak{g}$-module and let $\mathfrak{h}$ be a central ideal of $\mathfrak{g}$, such that $\mathfrak{h}\leq C_{\mathfrak{g}}(A)$. Suppose that $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g}/\mathfrak{h},A)\simeq H^{1}(\mathfrak{g},A)$.
Proof.
Since $A^{\mathfrak{h}}=A$, by Fact 7.4, the sequence $0\rightarrow H^{1}(\mathfrak{g}/\mathfrak{h},A)\overset{\inf}{\rightarrow}H^{1}(\mathfrak{g},A)\overset{\operatorname{res}}{\rightarrow}H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}$ is exact. Therefore, it suffices to prove that $H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}=0$.
Let $f\in\operatorname{Der}(\mathfrak{h},A)$; suppose that for all $\overline{x}\in\mathfrak{g}/\mathfrak{h}$, $(\overline{x}.f)]=[0]$, i.e., for all $y\in\mathfrak{h}$, we have $x.f(y)-f([x,y])=x.f(y)=y.a_{x}=0$ for some $a_{x}\in A$, since $\mathfrak{h}$ acts trivially and $\mathfrak{h}\leq Z(\mathfrak{g})$. In other words, for every $x\in\mathfrak{g}$, the equality $x\cdot f(y)=0$ holds, and thus $f(y)\in A^{\mathfrak{g}}=0$; finally, we obtain that $f=0$. ∎
Regarding the adaptation of the argument of section 3 and 4, the only minor variation concerns the abelian case for irreducible module.
Lemma 7.7.
Let $\mathfrak{g}$ be a finite-dimensional (in the usual linear sense) Lie algebra and $A$ be a finite-dimensional (in the usual linear sense) $\mathfrak{g}$-module. Suppose that $A$ is $\mathfrak{g}$-irreducible and $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g},A)=0$.
Proof.
We proceed by induction on the length of $A$ in the category of finite-dimensional $\mathfrak{g}$-modules. Since $A$ is finite-dimensional, the length is also finite. If $A=0$, we are done.
First, we way suppose that $\mathfrak{g}$ acts faithfully by Lemma 7.6 and so we have a embedding $\rho:\mathfrak{g}\hookrightarrow\operatorname{End}_{K}(A)$. Consider $R$ the subring of $\operatorname{End}_{K}(A)$ generated by $\mathfrak{g}$ and notice that $A$ is a simple $R$-module. By Schur’s Lemma, $C_{\operatorname{End}_{K}(A)}(R)$ is a skew-field containing $R$; up-to considering its center, we may suppose that it is commutative.
Now, let $\delta:\mathfrak{g}\rightarrow A$ be a $\mathfrak{g}$-derivation. Let $y\in\mathfrak{g}-\{0\}$; then $\rho(y)$ is invertible. Let $a=\rho(y)^{-1}\cdot\delta(y)$. For $x\in\mathfrak{g}$, by commutativity and the definition of a $\mathfrak{g}$-derivation, we have :
$\rho(x)\cdot a=\rho(x)\cdot\rho(y)^{-1}\cdot\delta(y)=\rho(y)^{-1}\cdot\rho(x)\cdot\delta(y)=\rho(y)^{-1}\cdot\rho(y)\cdot\delta(x)=\delta(x).$
Thus $\delta(x)=\delta_{a}(x)$, where $\delta_{a}(x)=\rho(x)\cdot a$ is the $\mathfrak{g}$-inner derivation associated with $a$. This shows that $H^{1}(\mathfrak{g},A)=0$. ∎
Now, we may generalize the statements of section 4 and section 5 to the case of Lie algebras/Lie rings definable in a finite-dimensional theory. Notice that the Frattini argument obtained for Lie algebras/Lie rings is stronger since the commutativity of the group structure is automatic. In particular, we have the following :
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie algebra/Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g},A)=0$.
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie algebra/Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A^{\mathfrak{g}}=0$. Then for all definable connected sub-modules $V\leq U$ contained in $A$, one has $(U/V)^{\mathfrak{g}}=0$.
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie algebra/Lie ring, and let $\mathfrak{c}$ be a definable connected Cartan subalgebra/subring of a definable connected ideal $\mathfrak{i}$. Then
    $\mathfrak{g}=\mathfrak{i}+N_{\mathfrak{g}}(\mathfrak{c}).$
Some applications of the Frattini argument will be given in an ongoing collaboration with Jules Tindzhogo Ntsiri, that deals with the structure of ranked soluble Lie rings through the analysis of their Cartan subrings (see [12]).
Regarding the adaptation of the content of section 6, we must clarify what semi-simplicity means for Lie algebras in positive characteristic. We will start with the algebra of endomorphisms of a finite-dimensional vector space, for which linear algebra provides a natural notion of semi-simple element. For the following developments, we refer to Chapter II of [6].
Let $V$ be a finite-dimensional vector space over a field $\mathbb{K}$ of characteristic $p>0$; consider the associative $\mathbb{K}$-algebra $\mathfrak{g}=\operatorname{End}(V)$. We aim to characterize semi-simple endomorphisms purely in terms of algebraic operations in $\mathfrak{g}$ without reference to the minimal polynomial. To this end, raising to the $p$-th power will play a key role; we begin with the following definition :
Definition 7.8.
An element $x$ of $\mathfrak{g}$ is algebraically semi-simple if
$x\in\langle\bigcup_{i\geq 1}x^{p^{i}}\rangle_{vect}.$
The following proposition shows that the two notions of semi-simplicity coincide :
Proposition 7.9.
[6, Proposition 3.3, chap. 2] An element $x$ of $\mathfrak{g}$ is algebraically semi-simple if and only if it induces a semi-simple endomorphism.
For a Lie algebra over a field of characteristic $p>0$, one needs an analogue of the $p$-power map; this leads to the notion of restricted Lie algebra :
Definition 7.10.
A Lie algebra $\mathfrak{g}$ over a field $\mathbb{K}$ of characteristic $p>0$ is restricted if there exists a map $p]:\mathfrak{g}\to\mathfrak{g}$, $x\mapsto x^{[p]}$, satisfying the following conditions :
  1. 1.
    $x^{[p]},y]=ad_{x}^{p}(y)$ for all $y\in\mathfrak{g}$.
  2. 2.
    $(\lambda x)^{[p]}=\lambda^{p}x^{[p]}$ for all $\lambda\in K$.
  3. 3.
    $(x+y)^{[p]}=x^{[p]}+y^{[p]}+\sum_{1\leq i\leq p-1}s_{i}(x,y)$, where $ad_{x\otimes X+y\otimes 1}^{p-1}(x\otimes 1)=\sum_{1\leq i\leq p-1}is_{i}(x,y)\otimes X^{i-1}$ in $\mathfrak{g}\otimes_{K}K[X]$, for all $x,y\in\mathfrak{g}$.
Moreover, if $\mathfrak{g}$ is abelian, then $s_{i}(x,y)=0$ for all $i\in\{1,\dots,p-1\}$ and all $x,y\in\mathfrak{g}$.
Note that an associative $\mathbb{K}$-algebra with its natural Lie algebra structure is restricted when $[p]$ is taken as the $p$-th power. In particular, the algebra of endomorphisms of a vector space over a field of positive characteristic and the universal enveloping algebra of a Lie algebra over a field of positive characteristic have a natural structure of restricted Lie algebras.
Definition 7.11.
Let $(\mathfrak{g},[p])$ be a restricted Lie algebra over a field of characteristic $p>0$. An element $x$ is semi-simple if
$x\in\langle\bigcup_{i\geq 1}x^{[p]^{i}}\rangle_{vect}.$
In a finite-dimensional theory, a definable connected restricted Lie algebra $(\mathfrak{t},[p])$ is a torus if it is abelian and $ker([p])=0$.
Remark 7.12.
In the linear finite-dimensional case, an element of a restricted Lie algebra over a perfect field of characteristic $p>0$ is semi-simple if and only if it belongs to an abelian subalgebra stable under $[p]$, such that $ker([p])=0$ in this subalgebra (see [6, Proposition 3.3 and Theorem 3.7]).
Now, we may use proposition 6.2 to prove the following :
Corollary 7.13.
In a finite-dimensional theory, let $\mathfrak{t}$ be a torus and $A$ a definable connected $\mathfrak{t}$-module. Assume that $A$ is abelian $p$-elementary and $\rho(t^{[p]})(x)=\rho(t)^{p}(x)$ for all $t\in\mathfrak{t}$ and $x\in A$, where $\rho$ is the morphism of representation induced by the action of $\mathfrak{t}$. Suppose that $A^{\mathfrak{t}}=0$. Then every definable connected irreducible (in the definable connected category) sub-module of $A$ admits a definable $\mathfrak{t}$-invariant complement.

### v2

@@
-Elements of Cohomology for model-theorically finite-dimensional groups and Lie algebras [^1] Keywords: nilpotent group, nilpotent Lie algebra, cohomology, finite-dimensional theory MSC 2020 : 17B56, 03C60
+On the cohomology of finite-dimensional nilpotent groups and Lie rings [^1] Keywords: nilpotent group, nilpotent Lie algebra, cohomology, finite-dimensional theory MSC 2020 : 17B56, 03C60
 Abstract
-We use the cohomology theory for groups invented by Hochshild and Serre to compute the first cohomology group for nilpotent groups that are definable in a finite-dimensional theory.
-Based on the established cohomological results, we derive some structural results : we prove a weak form of the Frattini argument for definable connected Cartan subgroups and we give a definable version of Maschke’s theorem in the case of a definable abelian $p^{\perp}$-group.
-The same results hold when one works with Lie algebras instead of groups.
+We give some vanishing results for the first classical cohomology group in the case of finite-dimensional nilpotent groups and Lie rings.
+Based on the established cohomological results, we derive some structural consequences : we prove a form of Frattini’s argument for definable connected Cartan subrings and we give a definable version of Maschke’s theorem for the action of a definable connected $p$-divisible abelian group.
 1 Introduction
 Under the impetus of Hochschild and Serre (see [8] and [9]), techniques from homological algebra have been successfully applied to the study of modules in the category of groups and of Lie algebras.
-For groups, cohomological methods play a prominent role in the classification of group extensions and had striking applications in number theory through Galois cohomology (see [10]).
-In the case of Lie algebras, a number of structural results on finite-dimensional (in the usual linear sense) solvable Lie algebras were obtained by Barnes by using cohomological tools (see [2], [3], and [4]).
-Therefore, it would be natural to generalize these methods to the context of groups/Lie algebras and associated modules when one works within the category of definable objects in a given first-order theory.
-We will consider the case of finite-dimensional theories (precise definitions will be given in the following).
-The existence of a finite dimension on the collection of interpretable sets, endowed in addition with good algebraic properties generalizing those of linear dimension, will enable us to have better control over constructions arising from homological algebra.
-Regarding the scope of our approach to cohomology in this context, we will use two main tools : 1) the exactness of the inflation-restriction sequence for $n=1$, which can be proved by using elementary computations on the cochains 2) the existence of long exact sequences of cohomology groups induced by short exact sequences of $G$-modules.
+For (finite) groups, cohomological methods play a prominent role in the classification of group extensions and had striking applications in number theory through Galois cohomology (see [10]).
+In the case of Lie algebras, a number of structural results for finite-dimensional (in the usual linear sense) solvable Lie algebras were obtained by Barnes by using cohomological tools (see [2] and [3]).
+The approach of Barnes is particulary relevant in our case; more precisely, in [2] he proved that the cohomology groups associated to a nilpotent Lie algebra of finite linear dimension vanishe if the submodule of invariants is trivial.
+The same result is true when one replaces nilpotent Lie algebras by nilpotent finite groups or algebraic groups, nevertheless, a general and uniform treatement for this kind of vanishing results seems to be lacking.
+Model-theory could provide this kind of general framework and we generalize Barnes’ methods to the context of groups and Lie rings (Lie algebras devoided of their vector space structure) equipped with some kind of dimension function on the definable subsets, which generalizes the properties of the linear dimension function and cardinality.
+From a model-theoretic point of view, we will consider groups and Lie rings that are definable in a finite-dimensional theory (see below for a precise definition).
+Finite groups, algebraic groups over an algebraically closed field, semi-algebraic groups over the reals are some striking examples of such "finite-dimensional" groups.
+On the other hand, finite-dimensional Lie algebras are the main instances of "finite-dimensional" Lie rings.
+Nevertheless, we decided to focus on the abstract setting of Lie rings since this kind of structures have recently captured the attention of the model-theoretic community (see [5] and [6] for example) and we hope what is presented here will contribute to this line of research.
+As an additional upshot, we revisite some standard algebraic results such as Frattini-like arguments and Maschke-like theorems.
+Before stating our results, we say some words regarding the scope of our approach to cohomology.
+We will use two main tools : 1) the exactness of the inflation-restriction sequence for $n=1$, which can be proved by using elementary computations on the cochains 2) the existence of "long" exact sequences of cohomology groups induced by short exact sequences of $G$-modules.
 In particular, the Hoshchild-Serre spectral sequence will play no role in our approach.
-We formulate our results for groups, but we could have opted for the point of view of Lie algebras (see the Appendix).
-The main results we prove are the following :
+Regarding cohomology, we prove the following :
+Theorem 1.1.
+  1. 1.
     In a finite-dimensional theory, let $G$ be a definable connected nilpotent group and $A$ be a definable connected $G$-module. Suppose that $A^{G}=0$. Then $H^{1}(G,A)=0$.
+  2. 2.
     In a finite-dimensional theory, let $G$ be a definable connected nilpotent group and $A$ be a definable connected $G$-module.
@@
 Then for all definable connected sub-modules $V\leq U$ contained in $A$, one has $(U/V)^{G}=0$.
-  • In a finite-dimensional theory, let $G$ be a definable connected group, and let $C$ be a definable connected Cartan subgroup of a definable connected normal subgroup $H$.
-Suppose that $G/N_{G}(C)$ is abelian.
-Then
-    $G=HN_{G}(C).$
-  • In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module $p$-elementary as an abelian group.
+We formulate our results for groups but we could have opted for the point of view of Lie rings. On the basis of our cohomological results, we give the two following algebraic applications:
+Theorem 1.2.
+  1. 1.
+    (definable Maschke’s theorem) In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module $p$-elementary as an abelian group.
 Suppose that $A^{T}=0$.
 Then $A=A_{1}\oplus\dots\oplus A_{r}$, where each $A_{i}$ is a definable connected $T$-irreducible sub-module of $A$.
-In the second section, we recall some basic results about the cohomology of groups.
-In the third section, we consider the case of a finite nilpotent group $G$ and a finite $G$-module and we give an elementary proof of the vanishing of the first cohomology group when the submodule of invariants is trivial.
-In the fourth section, we generalize the cohomological results for finite groups to the case of nilpotent groups definable in a finite-dimensional theory.
-In the fifth section, based on the established cohomological results, we derive a weak form of the Frattini argument for connected definable Cartan subgroups.
-In the sixth section, we prove a definable Maschke’s theorem for definable abelian groups without $p$-torsion acting on definable abelian $p$-elementary groups.
-Finally, in an appendix, we explain what happens to our results when we consider Lie algebras instead of groups.
-We thank Adrien Deloro for his many comments and suggestions.
-2 Preliminaries on the cohomology of groups
-We recall some basic facts regarding the cohomology of groups; for the sake of completeness, we give proofs when it is needed. For general references on this topic, we refer to [7] and to [10].
+  2. 2.
+    (Frattini’s argument for Lie rings) In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring, and let $\mathfrak{c}$ be a definable connected Cartan subring of a definable connected ideal $\mathfrak{i}$.Then
+    $\mathfrak{g}=\mathfrak{i}+N_{\mathfrak{g}}(\mathfrak{c}).$
+In the second section, we recall some basic results about the cohomology of groups and after reviewing the case of finite nilpotent groups, we give an elementary proof of the vanishing of the first cohomology group when the submodule of invariants is trivial for a definable connected nilpotent group in a finite-dimensional theory.
+In the third section, we give our first algebraic application and we prove a definable version of Maschke’s theorem for the action of a definable connected $p$-divisible abelian group.
+In the fourth section, we pass to the case of Lie rings and in particular, we give a version of Frattini’s argument for definable connected Cartan subrings.
+We thank Adrien Deloro for his many comments and suggestions. We are also grateful to Moreno Invitti for his key observation regarding long exact cohomological sequences.
+2 Cohomological results
+2.1 Preliminaries on the cohomology of groups
+We recall some basic facts regarding the cohomology of groups.
+Although the results presented in this subsection are fairly standard, we decided to provide complete elementary algebraic proofs that could easily generalize to the case of abstract Lie rings for which general cohomological results may fail.
+For general references on this topic, we refer to [8] and to [10].
 Definition 2.1.
@@
 Let $(G,A)$ be a $G$-module. We define a cochain complex as follows :
-  • $C^{0}(G,A)=A$
-  • For every integer $n\geq 1$, $C^{n}(G,A)$ denotes the abelian group of functions between $G^{n}$ and AA
+  • $C^{0}(G,A)=A$.
+  • For every integer $n\geq 1$, $C^{n}(G,A)$ denotes the abelian group of functions between $G^{n}$ and $A$.
 The differential $d_{n}$ between $C^{n}(G,A)$ and $C^{n+1}(G,A)$ is defined, for every map $f$ in $C^{n}(G,A)$, by:
@@
 $H^{n}(G,A)=\ker(d_{n})/\operatorname{im}(d_{n-1}).$
-The group $G$ acts on the cochain groups of the form $C^{n}(H,A)$, for $H$ a normal subgroup, in a way that is compatible with the differentials : let $f\in C^{n}(H,A)$ and $x\in G$, then $(x\cdot f)(h_{1},\dots,h_{n})=x\cdot f(h_{1}^{x},\dots,h_{n}^{x})$.
+The group $G$ acts on the cochain groups of the form $C^{n}(H,A)$, for $H$ a normal subgroup, in a way that is compatible with the differentials :
+let $f\in C^{n}(H,A)$ and $x\in G$, then $(x\cdot f)(h_{1},\dots,h_{n})=x\cdot f(h_{1}^{x},\dots,h_{n}^{x})$.
 For $n=0$ and $n=1$, it is possible to give a convenient algebraic description of the cohomology groups.
@@
 Remark 2.5.
-We observe that $G$ acts trivially on $H^{1}(G,A)$: let $f]\in H^{1}(G,A)$, where $f$ is a derivation, and let $x\in G$; then $(x\cdot f)(y)-f(y)=x\cdot f(y^{x})-f(y)=x\cdot(x^{-1}\cdot f(yx)+f(x^{-1}))-f(y)=f(yx)+f(1)-f(x)-f(y)=y\cdot f(x)+f(y)-f(x)-f(y)$ for all $y\in G$; in other words, $(x\cdot f)-f$ is an inner derivation and thus $x\cdot f]=[f]$.
-We now turn to the question of the relations between exact sequences of $G$-modules and cohomology.
+We observe that $G$ acts trivially on $H^{1}(G,A)$: let $f]\in H^{1}(G,A)$, where $f$ is a derivation, and let $x\in G$; then
+$(x\cdot f)(y)-f(y)=x\cdot f(y^{x})-f(y)$
+$=x\cdot(x^{-1}\cdot f(yx)+f(x^{-1}))-f(y)$
+$=f(yx)+f(1)-f(x)-f(y)=y\cdot f(x)+f(y)-f(x)-f(y)$
+$=y\cdot f(x)-f(x)$
+for all $y\in G$; in other words, $(x\cdot f)-f$ is an inner derivation and thus $x\cdot f]=[f]$.
+We now turn to the question of the relations between exact sequences of $G$-modules and long exact sequence of cohomology groups.
 Fact 2.6.
-Every short exact sequence of $G$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces a short exact sequence of complexes $0\rightarrow C(G,A)\overset{i^{*}}{\rightarrow}C(G,B)\overset{p^{*}}{\rightarrow}C(G,C)\rightarrow 0$, where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$.
+Every short exact sequence of $G$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces the following short exact sequence :
+$0\rightarrow C^{1}(G,A)\overset{i^{*}}{\rightarrow}C^{1}(G,B)\overset{p^{*}}{\rightarrow}C^{1}(G,C),$
+where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$. Moreover, $i^{*}$ and $p^{*}$ commute with the differential.
 Proof.
-Since $i$ and $p$ commute with the action of $G$, the formula for the differential allows us to show that $i^{*},p^{*}$ commute with the differential. Fix an integer $n$.
-Let us show that $i^{*}$ is injective.
-Indeed, for $\phi\in C^{n}(G,A)$, $i^{*}(\phi)=0$ iff $i(\phi(\overline{x}))=0$ for all $\overline{x}\in G^{n}$ iff $\phi(\overline{x})=0$ for all $\overline{x}\in G^{n}$ by injectivity of $i$.
+Since $i$ and $p$ commute with the action of $G$, the formula for the differential allows us to show that $i^{*},p^{*}$ commute with it.
+Let us show that $i^{*}$ is injective. Indeed, for $\phi\in C^{1}(G,A)$, $i^{*}(\phi)=0$ iff $i(\phi(x))=0$ for all $x\in G$ iff $\phi(x)=0$ for all $x\in G$ by injectivity of $i$.
 Now, we turn to the equality $\operatorname{im}(i^{*})=\ker(p^{*})$.
-Let $\phi\in\operatorname{im}(i^{*})\subseteq C^{n}(G,B)$; then there exists $\psi\in C^{n}(G,A)$, such that $\phi=i\circ\psi$ and hence $p^{*}(\phi)=p\circ i\circ\psi$.
-Consequently, for every $\overline{x}\in G^{n}$, $p(i(\psi(\overline{x}))=0$ since $\operatorname{im}(i)\subseteq\ker(p)$, and thus $\phi\in\ker(p^{*})$.
-Conversely, let $\phi\in\ker(p^{*})$; then $p(\phi(\overline{x}))=0$ for every $\overline{x}\in G^{n}$.
-Thus $\phi(\overline{x})\in\operatorname{im}(i)$, since $\operatorname{im}(i)=\ker(p)$.
-Therefore for every $\overline{x}$ there exists an unique $y_{\overline{x}}\in A$, such that $i(y_{\overline{x}})=\phi(\overline{x})$ by injectivity of $i$.
-We can define an element of $C^{n}(G,A)$ by $\overline{x}\mapsto y_{\overline{x}}$.
-Now, we prove the surjectivity of $p^{*}$.
-Let $s:C\rightarrow B$ be a set-theoretic section, i.e., $p\circ s=Id_{C}$.
-Let $\phi\in C^{n}(G,C)$; for $\overline{x}\in G$, consider $s(\phi(\overline{x}))$.
-Therefore, it suffices to define a cochain $\psi\in C^{n}(G,B)$ by $\psi(\overline{x})=s(\phi(\overline{x}))$ to get a map such that $p^{*}(\psi)=\phi$.
+Let $\phi\in\operatorname{im}(i^{*})\subseteq C^{1}(G,B)$; then there exists $\psi\in C^{1}(G,A)$, such that $\phi=i\circ\psi$ and hence $p^{*}(\phi)=p\circ i\circ\psi$.
+Consequently, for every $x\in G$, $p(i(\psi(x))=0$ since $\operatorname{im}(i)\subseteq\ker(p)$, and thus $\phi\in\ker(p^{*})$.
+Conversely, let $\phi\in\ker(p^{*})$; then $p(\phi(x))=0$ for every $x\in G$.
+Thus $\phi(x)\in\operatorname{im}(i)$, since $\operatorname{im}(i)=\ker(p)$.
+Therefore for every $x$ there exists an unique $y_{x}\in A$, such that $i(y_{x})=\phi(x)$ by injectivity of $i$.
+Now, it suffices to consider the following element of $C^{1}(G,A)$ : $\psi:x\mapsto y_{x}$.
 ∎
-Recall now an important property of the sequence of cohomology groups of a complex: to a short exact sequence of complexes, one can canonically associate a long exact sequence of cohomology groups.
-Fact 2.7.
-[14, Theorem 1.3.1], [9, Proposition 6.9]
-Let $0\rightarrow C^{\prime}\rightarrow C\rightarrow C^{\prime\prime}\rightarrow 0$ be a short exact sequence of cochain complexes.
-Then there exists a long exact sequence of cohomology groups of the following form :
-$\dots\rightarrow H^{n}(C^{\prime})\rightarrow H^{n}(C)\rightarrow H^{n}(C^{\prime\prime})$
-$\overset{\delta_{n}}{\rightarrow}H^{n+1}(C^{\prime})\rightarrow H^{n+1}(C)\rightarrow\dots$
-The morphisms $(\delta_{n})_{n\in\mathbb{Z}}$ are called connecting morphisms.
-As a consequence, an exact sequence of $G$-modules induces a long exact sequence of cohomology groups.
+Corollary 2.7.
+Every short exact sequence of $G$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces the following short exact sequence :
+$0\rightarrow Der(G,A)\overset{i^{*}}{\rightarrow}Der(G,B)\overset{p^{*}}{\rightarrow}Der(G,C),$
+where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$.
+Proof.
+Since $i$ and $p$ are $G$-morphisms, the sequence is well-defined.
+Moreover, let $\phi\in\ker(p^{*})$ and consider the application $\psi:G\rightarrow A,x\mapsto y_{x}$, where $i(y_{x})=\phi(x)$.
+Then, for all $x,y\in G$, we have $i(x\cdot\psi(y)+\psi(x))=x\cdot i(\psi(y))+i(\psi(x))=x\cdot\phi(y)+\phi(x)=\phi(xy)=i(\psi(xy))$, so, by injectivity of $i$, $\psi(xy)=x\cdot\psi(y)+\psi(x)$.
+∎
+One important feature of cohomology is the possibility of getting a "long" exact sequence of cohomology groups from a short exact sequence of $G$-modules.
+Fact 2.8.
+Let $0\rightarrow A_{0}\rightarrow A_{1}\rightarrow A_{3}\rightarrow 0$ be a short exact sequence of $G$-modules. Then there exists a "long" exact sequence of cohomology groups of the following form :
+$A_{0}^{G}\rightarrow A_{1}^{G}\rightarrow A_{2}^{G}$
+$\overset{\delta_{0}}{\rightarrow}H^{1}(G,A_{0})\rightarrow H^{1}(G,A_{1})\rightarrow\ H^{1}(G,A_{2})$
+The morphism $(\delta_{0})$ is called a connecting morphism.
+Proof.
+It suffices to apply the Snake Lemma to the following commutative diagram, where $\phi_{i}:A_{i}\rightarrow Der(G,A_{i}),a_{i}\mapsto(g\mapsto g\cdot a_{i}-a_{i}),i=0,1,2$ :
+[…]
 We now introduce the Hochschild-Serre inflation-restriction sequence.
@@
   • We define the inflation map $\inf:C^{n}(G/H,A^{H})\longrightarrow C^{n}(G,A),f\mapsto f\circ\pi$, where $\pi$ denotes the canonical projection between $G$ and $G/H$.
-Since the inflation and restriction morphisms commute with the differentials, we obtain a sequence of complexes.
-In particular, we deduce a sequence $H^{1}(G/H,A^{H})\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)$.
+Since the inflation and restriction morphisms commute with the differentials, we can pass to the cohomology groups and we get a sequence $H^{1}(G/H,A^{H})\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)$.
 In fact, this sequence is exact, and it can be proved by an elementary algebraic argument.
-Fact 2.9.
+Fact 2.10.
 [10, Proposition 4, §6, chap. VII])
@@
 We can refine the image of the restriction morphism.
-Fact 2.10.
+Fact 2.11.
 $H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)^{G/H}$.
@@
 Proof.
-Since $A^{H}=A$, by fact 2.9 and fact 2.10, the sequence $0\rightarrow H^{1}(G/H,A)\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)^{G/H}$ is exact.
+Since $A^{H}=A$, by Fact 2.10 and Fact 2.11, the sequence $0\rightarrow H^{1}(G/H,A)\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)^{G/H}$ is exact.
 Therefore, it suffices to prove that
@@
 ∎
-3 The case of finite groups
+2.2 The case of finite groups
 In this section, we give a quite elementary proof of the vanishing of the first cohomology group when $A^{G}=0$ for a finite nilpotent group $G$ and a finite $G$-module AA.
@@
 Let us recall a basic but fundamental lemma in representation theory : Schur’s Lemma.
-Fact 3.1.
+Fact 2.13.
 Let $A$ be a simple $R$-module, i.e $A$ has no non-trivial proper submodule. Then $C_{\operatorname{End}(A)}(R)$ is a skew-field.
@@
 Let $a=(\rho(y)-Id)^{-1}\cdot f(y)$.
-For $x\in G$, by commutativity of $G$ and the definition of a derivation, we have $(\rho(y)-Id)\cdot f(x)=\rho(x)\cdot f(y)-f(y)$, thus, $f(x)=(\rho(y)-Id)^{-1}\cdot\rho(x)\cdot f(y)-(\rho(y)-Id)^{-1}\cdot f(y)=\rho(x)\cdot a-a$ by commutativity again.
-This shows that $H^{1}(G,A)=0$.
+For $x\in G$, by commutativity of $G$ and the definition of a derivation, we have $(\rho(y)-Id)\cdot f(x)=\rho(x)\cdot f(y)-f(y)$, thus,
+    $f(x)=(\rho(y)-Id)^{-1}\cdot\rho(x)\cdot f(y)-(\rho(y)-Id)^{-1}\cdot f(y)=\rho(x)\cdot a-a$
+    by commutativity again. This shows that $H^{1}(G,A)=0$.
   • $\lg(A)>1$.
@@
 The long exact cohomology sequence will now come into play.
-Indeed, we have a short exact sequence of $G$-modules $0\rightarrow B\rightarrow A\rightarrow A/B\rightarrow 0$ which, by virtue of Fact 2.7, induces the following long exact cohomology sequence :
+Indeed, we have a short exact sequence of $G$-modules $0\rightarrow B\rightarrow A\rightarrow A/B\rightarrow 0$ which, by virtue of Fact 2.8, induces the following long exact cohomology sequence :
     $0\rightarrow B^{G}\rightarrow A^{G}\rightarrow(A/B)^{G}$
-    $\overset{\delta_{0}}{\rightarrow}H^{1}(G,B)\rightarrow H^{1}(G,A)\rightarrow H^{1}(G,A/B)\rightarrow\dots$
-    By hypothesis, $A^{G}=B^{G}=0$ and by induction hypothesis, $H^{1}(G,B)=0$, so the exactness of the sequence then gives us $(A/B)^{G}=0$; we may therefore apply once more the induction hypothesis and use the exactness of the long cohomology sequence to establish $H^{1}(G,A)=0$.
+    $\overset{\delta_{0}}{\rightarrow}H^{1}(G,B)\rightarrow H^{1}(G,A)\rightarrow H^{1}(G,A/B)$
+    By hypothesis, $A^{G}=B^{G}=0$ and by induction hypothesis, $H^{1}(G,B)=0$, so the exactness of the sequence then gives us that $(A/B)^{G}=0$; we may therefore apply once more the induction hypothesis and use the exactness of the long cohomology sequence to establish that $H^{1}(G,A)=0$.
 Corollary 2.15.
@@
 By irreducibility of $A$, we are reduced to two cases.
-  • $A^{Z(G)}=A$. Lemma 2.11 gives $H^{1}(Z(G),A)^{G/Z(G)}=0$; so, the inflation-restriction exact sequence entails $H^{1}(G/Z(G),A)\simeq H^{1}(G,A)$ and we are done by induction.
+  • $A^{Z(G)}=A$. Then Lemma 2.12 gives $H^{1}(Z(G),A)^{G/Z(G)}=0$; so, the inflation-restriction exact sequence entails $H^{1}(G/Z(G),A)\simeq H^{1}(G,A)$ and we are done by induction.
   • $A^{Z(G)}=0$. Since $H^{1}(Z(G),A)=0$ (abelian case) and $H^{1}(G/Z(G),0)=0$, we have $H^{1}(G,A)=0$ by considering again the inflation-restriction exact sequence.
@@
 To extend the proof to the category of definable $G$-modules in a first order theory, we need two main ingredients : 1) a notion of irreducible module that allows inductive reasoning on the lenght 2) an appropriate version of Schur’s Lemma regarding this irreducible category.
-In the next section, we will see that finite-dimensional theories satisfy these requirements.
-4 Finite-dimensional nilpotent groups
+In the next subsection, we will see that finite-dimensional theories satisfy these requirements.
+2.3 Finite-dimensional nilpotent groups
 We will study group cohomology inside the framework of finite-dimensional theories.
@@
   • (Fibration) If $f:X\longrightarrow Y$ is an interpretable map, such that $\dim(f^{-1}(y))\geq d$ for all $y\in Y$, then $\dim(X)\geq\dim(Y)+d$.
+  • (Low fibration) If $f:X\longrightarrow Y$ is an interpretable map, such that $\dim(f^{-1}(y))\leq d$ for all $y\in Y$, then $\dim(X)\leq\dim(Y)+d$.
 The dimension extends to type-definable sets, i.e., bounded intersections of definable sets, by taking the infimum of the dimensions of the definable sets which appear in the intersection.
-A strongly minimal theory, the theory of a group of finite Morley rank and of an o-minimal structure are examples of finite-dimensional theories (for o-minimal structures, we must restrict to the collection of definable sets and interpretable group quotients).
+The theory of a group of finite Morley rank and of an o-minimal structure are examples of finite-dimensional theories (for o-minimal structures, we must restrict to the collection of definable sets and interpretable group quotients).
 This model-theoretic framework is suitable for inductive reasoning when working with groups.
-Fact 4.2.
+Fact 2.18.
 [13, Proposition 2.3] Let $H\leq G$ be two type-definable groups in a finite-dimensional theory. If $H$ has infinite index in $G$, then $\dim(H)<\dim(G)$.
@@
 Another remarkable property of finite-dimensional theories concerns the possibility of obtaining a definable version of Schur’s Lemma.
-Fact 4.3.
+Fact 2.19.
 (adapted from [4]) Let $T$ be a finite-dimensional theory.
@@
 We denote by $\rho:G\rightarrow\operatorname{Aut}_{def}(A)\subseteq\operatorname{End}_{def}(A)$ the action.
-By Fact 4.3, we can linearize : there exists a definable skew-field $\mathbb{K}$ such that $\rho(G)\leq\mathbb{K}$.
+By Fact 2.19, we can linearize : there exists a definable skew-field $\mathbb{K}$ such that $\rho(G)\leq\mathbb{K}$.
 Up to considering $Z(\mathbb{K})$, which is also definable infinite, we may assume that $\mathbb{K}$ is commutative.
@@
 Any definable group admits a connected component, i.e., a smallest definable subgroup of finite index.
-A strongly minimal theory, the theory of an o-minimal structure, the theory of a group of finite Morley rank belong to this subclass.
+The theory of an o-minimal structure and the theory of a group of finite Morley rank belong to this subclass.
 In a finite-dimensional theory, a definable irreducible $G$-module has no proper infinite definable submodule.
@@
 In fact, $A^{Z(G)^{\circ}}=0$ since $A^{G}=0$ and $G$ is connected.
-Notice that $H^{1}(G/Z(G)^{\circ},A^{Z(G)^{\circ}})=0=H^{1}(Z(G)^{\circ},A)$ by Lemma 4.4 and we are done by fact 2.9.
+Notice that $H^{1}(G/Z(G)^{\circ},A^{Z(G)^{\circ}})=0=H^{1}(Z(G)^{\circ},A)$ by Lemma 2.20 and we are done by Fact 2.10.
 Now, we proceed by induction on the length of $A$ as a $G$-module exactly in the same way as in the proof of Lemma 2.20. ∎
@@
 $0\rightarrow V^{G}\rightarrow U^{G}\rightarrow(U/V)^{G}$
-$\overset{\delta_{0}}{\rightarrow}H^{1}(G,V)\dots$
+$\overset{\delta_{0}}{\rightarrow}H^{1}(G,V)$
 Theorem 2.22 allows us to conclude. ∎
-5 A weak Frattini argument for definable connected Cartan subgroups
-In the theory of finite groups, $p$-Sylow subgroups play a fundamental role and often allow to obtain structural results about the ambient group.
-In particular, the Frattini argument states that, for a finite group $G$, a normal subgroup $H$, and a $p$-Sylow subgroup $S$ of $H$, we have the following :
-$G=HN_{G}(S),$
-where $N_{G}(S)$ denotes the normalizer of $S$ in $G$.
-In fact, the image of $S$ under an inner automorphism $inn(g)$ remains a $p$-Sylow subgroup of $H$; by the conjugacy of $p$-Sylow subgroups in $H$, there exists $h\in H$, such that $S^{h}=S^{g}$, that is, such that $h^{-1}g\in N_{G}(S)$.
-Outside the world of finite groups, Cartan subgroups, i.e., nilpotent self-normalizing subgroups can be seen as a good proxy for $p$-Sylow subgroups.
-In general, the existence and the conjugation of Cartan subgroups are not guaranteed.
-Nevertheless, in a finite-dimensional theory, if there exist definable connected Cartan subgroups, the conclusion of the Frattini argument may still be true without any conjugation assumption.
-We follow the cohomological treatment for finite-dimensional Lie algebras given by Barnes in [4].
-Definition 5.1.
-In a finite-dimensional theory, let $G$ be a definable connected group and let $C$ be a definable connected nilpotent subgroup such that
-$N_{G}(C)=C.$
-We then say that $C$ is a definable connected Cartan subgroup.
-Proposition 5.2.
-In a finite-dimensional theory, let $G$ be a definable connected group, and let $C$ be a definable connected Cartan subgroup of a definable connected normal subgroup $H$.
-Suppose that $G/N_{G}(C)$ is abelian.
-Then
-$G=HN_{G}(C).$
-Proof.
-We set $N=N_{G}(C)$ and we consider the definable connected $C$-module $A=G/N$.
-Let us prove that $A^{C}=0$.
-Indeed, let $\overline{a}\in A$ be such that $c\cdot\overline{a}=\overline{a}$ for all $c\in C$; then $a^{-1}a^{c}\in H\cap N$.
-But $H$ is normal subgroup and $N_{H}(C)=C$, so $a^{-1}a^{c}\in C$ for all $c$, hence $a\in N_{G}(C)$.
-Now, consider the $C$-module
-$(G/N)/((HN)/N)\simeq G/(HN).$
-If this module were nontrivial, Corollary 4.7 would imply that $C$ acts non-trivially, which contradicts the fact that $H$ contains $C$. Therefore, we conclude that
-$G=HN.\qed$
 3 Definable Maschke’s theorem
-We will provide another application of the cohomological study of model-theorically finite-dimensional nilpotent groups by proving a definable version of Maschke’s theorem.
+In this section, we provide an application of the cohomological study of finite-dimensional nilpotent groups by proving a definable version of Maschke’s theorem.
 We will closely follow the argument for the case of ranked groups given by Tindzhogo Ntsiri in [11].
+Nevertheless, even for ranked groups, the result presented here is new since we can weaken the hypothesis "decent torus" and replace it by "definable connected abelian p-divisible".
 We will need the following definability lemma :
-Fact 6.1.
+Fact 3.1.
 ([13, Proposition 3.6], extracted from [4]) In a finite-dimensional theory, let $\mathbb{K}$ be a skew-field such that :
@@
 Proposition 3.2.
-In a finite-dimensional theory, let $V$ be a definable connected $p$-elementary abelian group and let $R$ be an invariant commutative ring of definable endomorphisms of $V$ generated by an infinite definable set $X$.
+In a finite-dimensional theory, let $V$ be a definable connected $p$-elementary abelian group and let $R$ be an unbounded invariant commutative ring of definable endomorphisms of $V$.
 Suppose that $R^{p}=R$.
-Let $W$ be a definable connected irreducible (in the definable category) sub-$R$-module, such that $V/W$ is also $R$-irreducible (in the definable category).
+Let $W$ be a definable connected irreducible sub-$R$-module, such that $V/W$ is also $R$-irreducible.
 Then $W$ admits a definable $R$-invariant complement.
 Proof.
-First, notice that $R$ is unbounded.
-By considering the action of $R$ on $W$ and $V/W$ and by applying Fact 4.3, the rings $C_{\operatorname{End}_{def}(W)}(R)=K_{W}$ and $C_{\operatorname{End}_{def}(V/W)}(R)=K_{V/W}$ are in fact definable skew-fields.
+By considering the action of $R$ on $W$ and $V/W$ and by applying Fact 2.19, the rings $C_{\operatorname{End}_{def}(W)}(R)=K_{W}$ and $C_{\operatorname{End}_{def}(V/W)}(R)=K_{V/W}$ are in fact definable skew-fields.
 Therefore $\operatorname{ann}_{R}(W)$ and $\operatorname{ann}_{R}(V/W)$ are prime ideals.
@@
 Indeed, let $0\neq r\in\operatorname{ann}_{R}(W)-\operatorname{ann}(V/W)$.
-The definable subgroups $\operatorname{im}(r)$ and $W\subseteq\ker(r)$ are $R$-invariant since $R$ is commutative.
-Then $\ker(r)/W$ and $(\operatorname{im}(r)+W)/W$ are $R$-invariant, and by irreducibility of $V/W$, we have $V=\operatorname{im}(r)+W$.
-Moreover, $\dim(V)=\dim(\operatorname{im}(r))+\dim(\ker(r))=\dim(\operatorname{im}(r))+\dim(\ker(r))-\dim(W\cap\operatorname{im}(r))$; thus $\dim(W\cap\operatorname{im}(r))=0$ and $\operatorname{im}(r)\cap W$ is trivial by irreducibility of $W$.
-Conversely, let $0\neq r\in\operatorname{ann}_{R}(V/W)-\operatorname{ann}_{R}(W)$. Then $\operatorname{im}(r)=W$ and $\ker(r)\cap W$ is trivial by irreducibility of $W$. Therefore $V=W\oplus\ker(r)$.
+The definable connected subgroups $\operatorname{im}(r)^{\circ}$ and $W\subseteq\ker(r)^{\circ}$ are $R$-invariant since $R$ is commutative.
+Then $\ker(r)^{\circ}/W$ and $(\operatorname{im}(r)^{\circ}+W)/W$ are $R$-invariant, and by irreducibility of $V/W$, we have $V=\operatorname{im}(r)^{\circ}+W$.
+Moreover, $\dim(V)=\dim(\operatorname{im}(r)^{\circ})+\dim(\ker(r)^{\circ})=\dim(\operatorname{im}(r)^{\circ})+\dim(\ker(r)^{\circ})-\dim(W\cap\operatorname{im}(r)^{\circ})$; thus $\dim(W\cap\operatorname{im}(r)^{\circ})=0$ and $\operatorname{im}(r)^{\circ}\cap W$ is finite.
+By linearizing again, $\operatorname{im}(r)^{\circ}\cap W$ is in fact trivial.
+Conversely, let $0\neq r\in\operatorname{ann}_{R}(V/W)-\operatorname{ann}_{R}(W)$.
+Then $\operatorname{im}(r)^{\circ}=W$ by hypothesis and irreducibility of $W$.
+But $\ker(r)^{\circ}\cap W$ is finite by irreducibility of $W$ and it is in fact trivial by linearizing again.
+Therefore $V=W\oplus\ker(r)^{\circ}$.
 Now, let $x\in\operatorname{ann}_{R}(W)=\operatorname{ann}_{R}(V/W)=I$; since $R=R^{p}$, there exists $y\in R$ such that $x=y^{p}$.
@@
 Let $Z\subseteq K$ be a type-definable subset.
-For $0\neq v\in V$, the evaluation map $r\mapsto r\cdot v$ is injective since $K$ acts by automorphisms; thus $\dim(Z)=\dim(Z\cdot v)\leq\dim(V)$.
+For $0\neq v\in V$, the evaluation map $K\rightarrow V,k\mapsto r\cdot v$ is injective since $K$ acts by automorphisms; thus $\dim(Z)=\dim(Z\cdot v)\leq\dim(V)$.
 By Lemma 3.1, $K$ is in fact a definable field.
@@
 Since $T$ is abelian and $T^{p}=T$, we have $R^{p}=R$.
-Moreover, $W$ and $V/W$ are $R$-irreducible (in the definable category) : let $W_{1}$ be a finite $R$-sub-module of $W$; in particular, this is a $T$-module, but, by connectedness of $T$, the action is trivial and $W_{1}=0$ since $V^{T}=0$; the proof for $V/W$ is similar since $(V/W)^{T}=0$.
+Moreover, $W$ and $V/W$ are $R$-irreducible.
 Then, we are done by Proposition 3.2.
@@
 ∎
-7 Appendix : Lie algebras
-This appendix deals with the case of Lie algebras. Indeed, in [8], Hochshild and Serre defines a cohomology theory for Lie algebras that parallels the theory for groups.
+4 The case of Lie rings
+4.1 Cohomological results
+In [9], Hochshild and Serre defines a cohomology theory for Lie algebras that parallels the theory for groups.
+The definition can easily be adapted to the more general context of Lie rings; nevertheless, general cohomological results for Lie algebras may fail when passing to Lie rings.
+Recall that a Lie ring $(\mathfrak{g},+,[,])$ is an abelian group equipped with a bi-additive and anti-symmetric function, $[,]$, that satisfies the Jacobi identity.
 Definition 4.1.
-A $\mathfrak{g}$-module is the data of ($\mathfrak{g}$,AA), where $\mathfrak{g}$ is a Lie algebra over a field $K$ and $A$ a $K$-vector space, equipped with a Lie algebra action $\cdot"$ of $\mathfrak{g}$ on $A$.
-Equivalently, this amounts to giving a Lie algebra homomorphism $\rho$ between $\mathfrak{g}$ and $\operatorname{End}_{K}(A)$.
+A $\mathfrak{g}$-module is the data of ($\mathfrak{g}$,AA), where $\mathfrak{g}$ is a Lie ring and $A$ is an abelian group, equipped with a Lie ring action $\cdot"$ of $\mathfrak{g}$ on $A$.
+Equivalently, this amounts to giving a Lie ring homomorphism $\rho$ between $\mathfrak{g}$ and $\operatorname{End}(A)$.
 Definition 4.2.
@@
   • $C^{0}(\mathfrak{g},A)=A$
-  • For every integer $n\geq 1$, $C^{n}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g}^{n},A):f\text{ is alternating}\}$, where $\operatorname{Hom}(\mathfrak{g}^{n},A)$ denotes the set of applications between $\mathfrak{g}^{n}$ and $A$, which are $K$-linear in each variable (we say also $K$-multilinear).
-Recall that a $K$-multilinear application $f$ is alternating if $f(x_{1},\dots,x_{n})=0$ whenever two arguments among the $x_{i}$ are equal.
+  • For every integer $n\geq 1$, $C^{n}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g}^{n},A):f\text{ is alternating}\}$, where $\operatorname{Hom}(\mathfrak{g}^{n},A)$ denotes the set of applications between $\mathfrak{g}^{n}$ and $A$, which are $\mathbb{Z}$-linear in each variable (we say also $\mathbb{Z}$-multilinear).
+Recall that a $\mathbb{Z}$-multilinear application $f$ is alternating if $f(x_{1},\dots,x_{n})=0$ whenever two arguments among the $x_{i}$ are equal.
 The differential $d_{n}$ between $C^{n}(\mathfrak{g},A)$ and $C^{n+1}(\mathfrak{g},A)$ is defined, for every map $f$ in $C^{n}(\mathfrak{g},A)$, by :
@@
 $+\sum_{1\leq s\leq n+1}(-1)^{s}g_{s}\cdot f(g_{1},\dots,\hat{g_{s}},\dots,g_{n+1})$
-The Lie algebra $\mathfrak{g}$ acts on the cochain groups in a way that is compatible with the differentials :
+The Lie ring $\mathfrak{g}$ acts on the cochain groups in a way that is compatible with the differentials :
 Let $f\in C^{n}(\mathfrak{g},A)$ and $x\in\mathfrak{g}$, we define $x\cdot f$ by the formula
@@
 For $n=0$ and $n=1$, it is also possible to give a convenient algebraic description of the cohomology groups.
-Fact 7.3.
+Fact 4.3.
   1. 1.
@@
     $\operatorname{IDer}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g},A):f(x)=x.a\penalty 10000\ \text{for some $a$ in $A$}\}.$
-All the others statements of section 2 are still true when passing to Lie algebras, and the proofs require minor adaptations. In particular, we have the following :
-Fact 7.4.
-  • Every short exact sequence of $\mathfrak{g}$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces a short exact sequence of complexes $0\rightarrow C(\mathfrak{g},A)\overset{i^{*}}{\rightarrow}C(\mathfrak{g},B)\overset{p^{*}}{\rightarrow}C(\mathfrak{g},C)\rightarrow 0$, where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$.
+All the statements that we proved regarding the cohomology of groups are still true when passing to Lie rings, and the proofs require minor adaptations. In particular, we have the following :
+Fact 4.4.
+  • Every short exact sequence of $\mathfrak{g}$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces a "long" exact sequence of cohomology groups of the following form :
+    $A^{\mathfrak{g}}\rightarrow B^{\mathfrak{g}}\rightarrow C^{\mathfrak{g}}$
+    $\overset{\delta_{0}}{\rightarrow}H^{1}(\mathfrak{g},A)\rightarrow H^{1}(\mathfrak{g},B)\rightarrow\ H^{1}(\mathfrak{g},C)$
   • The sequence $0\rightarrow H^{1}(\mathfrak{g}/\mathfrak{h},A^{\mathfrak{h}})\overset{\inf}{\rightarrow}H^{1}(\mathfrak{g},A)\overset{\operatorname{res}}{\rightarrow}H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}$ is exact.
-Remark 7.5.
-As pointed out by Moreno Invitti, we can obtain the first six terms of the long exact sequence of cohomology groups induced by a short exact sequence of $\mathfrak{g}$-modules when $\mathfrak{g}$ is an abstract Lie ring.
-Indeed, it suffices to apply the Snake Lemma to the commutative diagram whose lines are $A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ and $0\rightarrow\operatorname{Der}(\mathfrak{g},A)\overset{i^{*}}{\rightarrow}\operatorname{Der}(\mathfrak{g},B)\overset{p^{*}}{\rightarrow}\operatorname{Der}(\mathfrak{g},C)$.
-Since we do not use the long exact sequence of cohomology groups for higher values of $n$, all the results of this appendix are still true for Lie rings that are definable in a finite-dimensional theory.
-Some generalizations of the results of this paper are currently under investigation by Moreno Invitti.
 Lemma 4.5.
@@
 Proof.
-Since $A^{\mathfrak{h}}=A$, by Fact 7.4, the sequence $0\rightarrow H^{1}(\mathfrak{g}/\mathfrak{h},A)\overset{\inf}{\rightarrow}H^{1}(\mathfrak{g},A)\overset{\operatorname{res}}{\rightarrow}H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}$ is exact.
+Since $A^{\mathfrak{h}}=A$, by Fact 4.4, the sequence $0\rightarrow H^{1}(\mathfrak{g}/\mathfrak{h},A)\overset{\inf}{\rightarrow}H^{1}(\mathfrak{g},A)\overset{\operatorname{res}}{\rightarrow}H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}$ is exact.
 Therefore, it suffices to prove that $H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}=0$.
@@
 ∎
-Regarding the adaptation of the argument of section 3 and 4, the only minor variation concerns the abelian case for irreducible module.
+There is also a minor variation concerning the abelian case for irreducible module.
 Lemma 4.6.
-Let $\mathfrak{g}$ be a finite-dimensional (in the usual linear sense) Lie algebra and $A$ be a finite-dimensional (in the usual linear sense) $\mathfrak{g}$-module.
+In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring and $A$ be a definable connected $\mathfrak{g}$-module.
 Suppose that $A$ is $\mathfrak{g}$-irreducible and $A^{\mathfrak{g}}=0$.
@@
 Proof.
-We proceed by induction on the length of $A$ in the category of finite-dimensional $\mathfrak{g}$-modules. Since $A$ is finite-dimensional, the length is also finite. If $A=0$, we are done.
-First, we way suppose that $\mathfrak{g}$ acts faithfully by Lemma 7.6 and so we have a embedding $\rho:\mathfrak{g}\hookrightarrow\operatorname{End}_{K}(A)$.
-Consider $R$ the subring of $\operatorname{End}_{K}(A)$ generated by $\mathfrak{g}$ and notice that $A$ is a simple $R$-module.
-By Schur’s Lemma, $C_{\operatorname{End}_{K}(A)}(R)$ is a skew-field containing $R$; up-to considering its center, we may suppose that it is commutative.
+First, we way suppose that $\mathfrak{g}$ acts faithfully by Lemma 4.5 and so we have an embedding $\rho:\mathfrak{g}\hookrightarrow\operatorname{End}_{def}(A)$.
+Consider $R$ the subring of $\operatorname{End}_{def}(A)$ generated by $\mathfrak{g}$; by Fact 2.19 , $C_{\operatorname{End}_{def}(A)}(R)$ is a definable skew-field containing $R$; up-to considering its center, we may suppose that it is commutative.
 Now, let $\delta:\mathfrak{g}\rightarrow A$ be a $\mathfrak{g}$-derivation.
@@
 Thus $\delta(x)=\delta_{a}(x)$, where $\delta_{a}(x)=\rho(x)\cdot a$ is the $\mathfrak{g}$-inner derivation associated with $a$. This shows that $H^{1}(\mathfrak{g},A)=0$. ∎
-Now, we may generalize the statements of section 4 and section 5 to the case of Lie algebras/Lie rings definable in a finite-dimensional theory.
-Notice that the Frattini argument obtained for Lie algebras/Lie rings is stronger since the commutativity of the group structure is automatic.
-In particular, we have the following :
-  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie algebra/Lie ring and $A$ be a definable connected $\mathfrak{g}$-module.
+Finally, we have the following :
+Theorem 4.7.
+  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie ring and $A$ be a definable connected $\mathfrak{g}$-module.
 Suppose that $A^{\mathfrak{g}}=0$.
 Then $H^{1}(\mathfrak{g},A)=0$.
-  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie algebra/Lie ring and $A$ be a definable connected $\mathfrak{g}$-module.
+  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie ring and $A$ be a definable connected $\mathfrak{g}$-module.
 Suppose that $A^{\mathfrak{g}}=0$.
 Then for all definable connected sub-modules $V\leq U$ contained in $A$, one has $(U/V)^{\mathfrak{g}}=0$.
-  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie algebra/Lie ring, and let $\mathfrak{c}$ be a definable connected Cartan subalgebra/subring of a definable connected ideal $\mathfrak{i}$.
-Then
-    $\mathfrak{g}=\mathfrak{i}+N_{\mathfrak{g}}(\mathfrak{c}).$
-Some applications of the Frattini argument will be given in an ongoing collaboration with Jules Tindzhogo Ntsiri, that deals with the structure of ranked soluble Lie rings through the analysis of their Cartan subrings (see [12]).
-Regarding the adaptation of the content of section 6, we must clarify what semi-simplicity means for Lie algebras in positive characteristic.
-We will start with the algebra of endomorphisms of a finite-dimensional vector space, for which linear algebra provides a natural notion of semi-simple element.
+4.2 Definable Maschke’s theorem for restricted Lie algebras
+Regarding Maschke’s theorem, we must pass to Lie algebras since the abstract Lie ring structure seems to be insufficient to capture a good notion of $p$-divisibility/$p$-unipotency.
+First, we clarify what semi-simplicity means for Lie algebras in positive characteristic.
+We start with the algebra of endomorphisms of a finite-dimensional vector space, for which linear algebra provides a natural notion of semi-simple element.
 For the following developments, we refer to Chapter II of [7].
@@
 $x\in\langle\bigcup_{i\geq 1}x^{p^{i}}\rangle_{vect}.$
-The following proposition shows that the two notions of semi-simplicity coincide :
-Proposition 7.9.
-[6, Proposition 3.3, chap. 2] An element $x$ of $\mathfrak{g}$ is algebraically semi-simple if and only if it induces a semi-simple endomorphism.
+The following fact shows that the two notions of semi-simplicity coincide :
+Fact 4.9.
+[7, Proposition 3.3, chap. 2] An element $x$ of $\mathfrak{g}$ is algebraically semi-simple if and only if it induces a semi-simple endomorphism.
 For a Lie algebra over a field of characteristic $p>0$, one needs an analogue of the $p$-power map; this leads to the notion of restricted Lie algebra :
@@
 Remark 4.12.
-In the linear finite-dimensional case, an element of a restricted Lie algebra over a perfect field of characteristic $p>0$ is semi-simple if and only if it belongs to an abelian subalgebra stable under $[p]$, such that $ker([p])=0$ in this subalgebra (see [6, Proposition 3.3 and Theorem 3.7]).
-Now, we may use proposition 6.2 to prove the following :
+In the linear finite-dimensional case, an element of a restricted Lie algebra over a perfect field of characteristic $p>0$ is semi-simple if and only if it belongs to an abelian subalgebra stable under $[p]$, such that $ker([p])=0$ in this subalgebra (see [7, Proposition 3.3 and Theorem 3.7]).
+Now, we may use Proposition 3.2 combining with Theorem 4.7 to prove the following :
 Corollary 4.13.
-In a finite-dimensional theory, let $\mathfrak{t}$ be a torus and $A$ a definable connected $\mathfrak{t}$-module.
-Assume that $A$ is abelian $p$-elementary and $\rho(t^{[p]})(x)=\rho(t)^{p}(x)$ for all $t\in\mathfrak{t}$ and $x\in A$, where $\rho$ is the morphism of representation induced by the action of $\mathfrak{t}$.
-Suppose that $A^{\mathfrak{t}}=0$.
-Then every definable connected irreducible (in the definable connected category) sub-module of $A$ admits a definable $\mathfrak{t}$-invariant complement.
+In a finite-dimensional theory, let $\mathfrak{t}$ be a torus and $V$ a definable connected $\mathfrak{t}$-module.
+Assume that $V$ is abelian $p$-elementary and $\rho(t^{[p]})(x)=\rho(t)^{p}(x)$ for all $t\in\mathfrak{t}$ and $x\in V$, where $\rho$ is the morphism of representation induced by the action of $\mathfrak{t}$.
+Suppose that $V^{\mathfrak{t}}=0$.
+Then $V=V_{1}\oplus\dots\oplus V_{r}$, where each $V_{i}$ is a $\mathfrak{t}$-irreducible definable connected sub-module.
+4.3 Frattini’s argument for Lie rings
+In the theory of finite groups, $p$-Sylow subgroups play a fundamental role and often allow to obtain structural results about the ambient group.
+In particular, Frattini’s argument states that, for a finite group $G$, a normal subgroup $H$, and a $p$-Sylow subgroup $S$ of $H$, we have the following :
+$G=HN_{G}(S).$
+The argument is based on conjugacy results for $p$-Sylow and thus generalizes poorly to the setting of Lie rings : even when one has a suitable analogue of $p$-Sylow subgroups, the "inner" maps are not automorphisms of the Lie ring structure, and passing through the exponential map only partially remedies this defect.
+However, the conclusion of Frattini’s argument remains valid for finite-dimensional (in the usual linear sense) Lie algebras when $p$-Sylow subgroups are replaced by Cartan subalgebras.
+The proof relies on the cohomological study of nilpotent Lie algebras (see [3, Theorem 2.1]).
+Taking into account our cohomological results and following [3, Theorem 2.1], we are now able to prove an analogous result in the context of finite-dimensional theories.
+Definition 4.14.
+In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring and let $\mathfrak{c}$ be a definable connected nilpotent subring such that
+$N_{\mathfrak{g}}(\mathfrak{c})=\mathfrak{c}.$
+We then say that $\mathfrak{c}$ is a definable connected Cartan subring.
+Proposition 4.15.
+In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring, and let $\mathfrak{c}$ be a definable connected Cartan subring of a definable connected ideal $\mathfrak{i}$. Then
+$\mathfrak{g}=\mathfrak{i}+N_{\mathfrak{g}}(\mathfrak{c}).$
+Proof.
+We set $\mathfrak{n}=N_{\mathfrak{g}}(\mathfrak{c})$ and we consider the definable connected $\mathfrak{c}$-module $A=\mathfrak{g}/\mathfrak{n}$.
+Let us prove that $A^{\mathfrak{c}}=0$ for the adjoint action.
+Indeed, let $\overline{a}\in A$ be such that $c\cdot\overline{a}=\overline{0}$ for all $c\in\mathfrak{c}$; then $c,a]\in\mathfrak{i}\cap\mathfrak{n}$.
+But $\mathfrak{i}$ is an ideal and $N_{\mathfrak{i}}(\mathfrak{c})=\mathfrak{c}$, so $c,a]\in\mathfrak{c}$ for all $c$, hence $\overline{a}=\overline{0}$.
+Now, consider the $\mathfrak{c}$-module
+$(\mathfrak{g}/\mathfrak{n})/((\mathfrak{i}+\mathfrak{n})/\mathfrak{n})\simeq\mathfrak{g}/(\mathfrak{i}+\mathfrak{n}).$
+If $(\mathfrak{g}/\mathfrak{n})/((\mathfrak{i}+\mathfrak{n})/\mathfrak{n})\neq 0$, by Theorem 4.7, $\mathfrak{c}$ acts non-trivially, a contradiction since $\mathfrak{c}\leq\mathfrak{i}$ and $\mathfrak{i}$ is an ideal.
+Some applications of Frattini’s argument will be given in [12] (in collaboration with Tindzhogo Ntsiri).
+Notably, we prove that the Frattini subring is nilpotent in the context of ranked connected soluble Lie rings.
+A generalization of some parts of this paper is currently under investigation by Moreno Invitti.